OpenStudy (anonymous):

3p over/ 8 - 13 = -25 a. -32 b. -101 1/3 c. -27 d. -45

sammixboo (sammixboo):

So does your equation look like this: \[\frac{ 3p }{ 8 } - 13 = -25\]

OpenStudy (anonymous):

yes @sammixboo

OpenStudy (anonymous):

1) add 13 to both sides 2) multiply both sides by 8 3) divide both sides by 3. (instead of step 2 and step 3, you can just multiply by 8/3 it's all same thing. )

OpenStudy (anonymous):

Let him try, sammixboo.

sammixboo (sammixboo):

You always have to get the variable by itself first, so you have to first divide both sides by 3, and then multiply both sides by 8. The last step is to add 13 to both sides, and there is your answer....

sammixboo (sammixboo):

I never gave the answer...

OpenStudy (anonymous):

Nooo !! You are wrong....

OpenStudy (anonymous):

The steps are wrong and out of order

sammixboo (sammixboo):

Well that is how I learned it... Just saying.

OpenStudy (anonymous):

You have the correct stpes, but they are in the wrong order... also, if you divide by 3 and multiply times 8, you won't have 13 in the equation anymore.

OpenStudy (anonymous):

\(\LARGE\color{black}{ \bf \frac{3p}{8} -13=-25 \ }\) \(\LARGE\color{black}{ \bf \frac{3p}{8} =-12 \ }\) \(\LARGE\color{black}{ \bf 3p =-96 }\) \(\LARGE\color{black}{ \bf p =-32 }\)